derivation of (a-b)^2: derivation of (a-b)^2
Answer by stanbon(19673)

derivation of (a-b)^2
----------------
(a-b)(a-b)
= a(a-b) -b(a-b)
= a^2 -ab -ba + b^2
= a^2 -ab - ab + b^2
= a^2 -2ab + b^2
====================
Cheers,
Stan H.